OpenStudy (randomchild):

In a piece-wise function the limit from the left and right approaches infinity but the function is defined for a finite number at the limit. Is it more correct to put the value of the limit as the finite value or infinity?

OpenStudy (anonymous):

infinity. the finite value doesn't necessarily tell you about the limit

OpenStudy (anonymous):

for example \(f(x) = \frac{1}{x^2}, x \neq 0\) \(a, x = 0\) f(0) = a, but \[\lim_{x \rightarrow 0}f(x)=\infty\]